Flying into the wind, an aircraft made a 360 km trip in 2 hours. The return trip with a tail wind took 90 minutes. Find the speed of the plane and the wind.
Answer by Alan3354(31518)

Flying into the wind, an aircraft made a 360 km trip in 2 hours. The return trip with a tail wind took 90 minutes. Find the speed of the plane and the wind.
------------
360/2 = 180 km/hr
360/1.5 = 240 km/hr
The airspeed is the average = (180 + 240)/2 = 210 km/hr
The windspeed is the difference = 240 - 210 = 30 km/hr